Circular No. 7899

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMETS C/2000 H6, H7, J8, N4; C/2002 J3 (SOHO) Further to IAUC 7898, D. Hammer reports his measurements for ad-

ditional comets found by R. Kracht (2000 comets; C2 coronagraph) and by K. Černis (C/2002 J3; C3) in the SOHO archive. The reduced astrometry and orbital elements by B. G. Marsden appear on the MPECs cited below. Comets C/2000 H6, C/2000 H7, and C/2002 J3 are members of the Kreutz group; the other two belong to the Meyer group (cf. IAUC 7832).

Comet  2000 UT  α_{2000}  δ_{2000}  MPEC
C/2000 H6 Apr. 26.063  2^{h}18.^{m}6  +12^{o}14′  2002-H49
C/2000 H7  28.663  2 30.2  +13 11  2002-H49
C/2000 J8 May 6.226  2 58.0  +17 46  2002-H49
C/2000 N4 July 4.660  6 55.0  +23 50  2002-J09
Comet  2002 UT  α_{2000}  δ_{2000}  MPEC
C/2002 J3 May 13.446  3^{h}32.^{m}7  +14^{o}14′  2002-J46

COMET C/2002 J4 (NEAT) S. Pravdo, E. Helin, and K. Lawrence, Jet Propulsion Laboratory,

report the discovery by NEAT, on CCD images taken on May 13.5 UT with the Palomar 1.2-m Schmidt telescope, of a comet with nebulosity extending ∼ 5′′ toward the southwest. Following posting on the NEO Confirmation Page, additional observations obtained elsewhere on May 13.9 confirm the cometary nature: P. Kušnirák at Ondřejov (0.65-m reflector; m_{1} = 18.3, faint 7′′ coma slightly elongated to the southwest); M. Tichý and M. Kočer at Kleť (1.06-m KLENOT telescope; compact coma of diameter 16′′); and L. Kornoš at Modra (0.6-m reflector; m_{1} = 16.6, coma diameter ∼ 15′′). Additional observations made by LONEOS and LINEAR on May 4, 7, and 8 were then identified by the Minor Planet Center staff.

2002 UT  α_{2000}  δ_{2000}  m_{1}
May 13.45406  15^{h}49^{m}37.^{s}19  +0^{o}26′39′′.8  18.0

Astrometry and the following preliminary parabolic orbital elements appear on MPEC 2002-J47:

T = 2003 Oct. 1.301 TT  ω = 230.^{o}588
Ω = 70.992 }2000.0
q = 3.62793 AU  i = 46.333
2002 May 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green